# Electricity Demand and Grid Impacts of AI Data Centers: Challenges and Prospects

### v3

@@ I Introduction
In recent years, the accelerated advancement of generative artificial intelligence (AI), particularly large language models (LLMs) [1] such as GPT [2], LLaMA [3], and Gemini [4], has fueled explosive growth in the AI industry. This surge is driving the rapid expansion of data center infrastructure and imposing unprecedented pressure on the power grid due to the immense electricity demands of ultra-scale AI workloads. For instance, training GPT-3 is estimated to have consumed 1.29 GWh of electricity [5], whereas the electricity consumption for training the larger-scale GPT-4 rose dramatically to an estimated over 50 GWh [6, 7], equivalent to nearly 0.1% of New York City’s annual electricity use. According to a recent International Energy Agency (IEA) report [8], global data centers consumed around 415 TWh of electricity in 2024 (about 1.5% of total global demand), and their consumption is projected to more than double by 2030 to around 945 TWh, with AI identified as the primary driver of this growth.
@@
AI data centers are the computing facilities that are designed and optimized to execute large-scale AI workloads, such as the training and inference of LLMs, computer vision systems, and other compute-intensive AI applications. In contrast to traditional data centers, which primarily provide general-purpose information technology (IT) services, AI data centers are architected to deliver the extreme computational performance required for advanced AI workloads (see Table I for a detailed comparison). This capability is enabled by dense configurations of high-performance computing hardware, such as Graphics Processing Units (GPUs) [9] and Tensor Processing Units (TPUs) [10], together with high-efficiency cooling systems [11]. Modern hyperscale AI data centers typically operate with power demand exceeding 100 MW [12], and some new campuses are planned to scale to the gigawatt (GW) level [13]. As a result, AI data centers are emerging as prominent large electric loads in the power grid, with distinct characteristics and demand patterns, including:
@@
  • High power density: Intensive AI workloads require substantially more power than conventional IT services. For instance, a ChatGPT query is estimated to consume about 2.9 Wh, nearly ten times the 0.3 Wh of a regular Google search [14]. In contrast to traditional server racks typically operating at 7-10 kW, AI computing racks can reach power densities of 30-100+ kW per rack [15]. Such extreme power density places considerable strain on data center architecture, cooling systems, and power grids.
@@
  • Fast and large variability: Across the training, fine-tuning, and inference phases, AI workloads can be highly variable and bursty [16], with power demand fluctuating sharply over very short timescales and remaining difficult to forecast [17]. For example, large-scale GPU clusters can produce power fluctuations of hundreds of megawatts within only seconds [18, 19], bringing significant challenges for system balancing and reliable grid operation.
@@
  • Grid interface via power electronics: AI data centers interface with the power grid primarily through power electronic converters, which exhibit fundamentally different dynamic behaviors from conventional electromechanical loads, such as low inertia, fast response dynamics, and harmonic distortions [20]. At scale, power electronics-based AI computing loads can pose major threats to grid stability and cause severe power quality issues [21].
@@
  • Geographic concentration: AI data centers tend to cluster in regions with low electricity prices, ample land and water resources, and favorable policy environments. In the U.S., for example, fifteen states (most notably Virginia, Texas, and California) accounted for about 80% of total data center demand in 2023 [14]. Such concentration effect further intensifies local grid stress, and many regional grids cannot accommodate large-scale integration of data centers without substantial infrastructure upgrades [22].
@@
A growing number of recent articles and reports [23, 7, 8, 24, 14, 25] have drawn considerable attention to the surging electricity demand of AI data centers and their implications for the power sector. Prior works [11, 14] examine the energy use of AI data centers by analyzing the power consumption of key components, including AI computing, cooling systems, and auxiliary facilities, and present strategies for energy-efficient operation. Reference [26] proposes a benchmarking framework for quantifying the footprint of energy, water, and carbon emissions of LLM inference workloads. Recent study [16] analyzes the transient dynamic behaviors of AI computing power consumption, develops high-level mathematical models for AI data center loads, and discusses the resulting disruptions and opportunities for power systems. Case studies on several power grids are conducted in [22] to assess their capacity to accommodate projected five-year AI data center load growth, deriving implications for long-term grid planning and management. However, the existing literature has mainly focused on the AI data center itself and presented high-level, preliminary, and fragmented discussions, while a comprehensive and in-depth synthesis of AI data center load characteristics and their impacts on power grids, particularly the challenges from the grid’s perspective, remains lacking.
@@ II-A1 Computing
AI workloads, such as LLM training and inference, require high-performance computing clusters designed to execute large-scale computations efficiently. These clusters typically integrate multi-core Central Processing Units (CPUs) with dense configurations of GPUs and TPUs [10]. These processing units are housed in high-density server nodes and stacked in standardized computer racks, allowing scalable deployment, efficient power distribution, and optimized cooling. Compared with conventional data centers, AI-oriented computing facilities concentrate substantially greater processing power within each rack. While a standard rack typically uses 7-10 kilowatt (kW), an AI-capable rack can demand 30 kW to over 100 kW [14, 25], with an average of more than 60 kW in dedicated AI facilities [15]. For example, the per-rack power of NVIDIA’s GB200 NVL36 and NVL72 servers reaches about 66 kW and 120 kW [27], respectively. The substantially higher power densities pose critical challenges for power distribution and thermal management.
@@ II-A2 Storage
To sustain intense compute capabilities, scalable and high‑performance data storage is necessitated to handle massive training datasets, intermediate model checkpoints, and inference data. Tiered storage components include: (i) the hot tier, such as Non-Volatile Memory Express (NVMe) Solid State Drives (SSDs) arrays [28], provides ultra-low latency and high bandwidth to store active data for fast processing, (ii) the warm tier, such as high-capacity hard disk drives (HDDs), is cost-effective for storing massive, less frequently accessed data, and (iii) the code tier stores data in tape libraries or object storage systems for long-term archiving with less access speed. In addition, distributed file systems, such as Lustre and Ceph [29], allow large‑scale parallel reading and writing, while object storage is used for archiving large datasets with long durability and redundancy [30]. Unlike conventional data centers, which balance capacity, cost, and speed, AI-optimized storage prioritizes maximum throughput and parallel access to prevent bottlenecks that idle expensive GPU clusters.
@@ II-A3 Networking
High‑speed networking is critical to ensure that massive data flow efficiently between storage systems and compute nodes during large‑scale AI workloads. The networking system in an AI data center includes both intra-node interconnects, such as NVIDIA NVLink [31] for high-bandwidth GPU-to-GPU communication within a server, and inter-node fabrics, such as InfiniBand [32] or Ethernet with RDMA over Converged Ethernet (RoCE), for low-latency, high-throughput communication between servers and racks. Specialized network topologies, including fat‑tree and dragonfly architectures [33], are adopted to optimize GPU‑to‑GPU communication in distributed training environments. In contrast, conventional data centers are usually designed for more heterogeneous workloads and general-purpose traffic patterns, and do not require such high bandwidth or low latency.
@@ II-B1 Grid Power Supply
AI data centers are typically connected to the medium- or high-voltage distribution grid, while hyperscale facilities may connect directly to the transmission grid as their primary power source. Depending on the local grid infrastructure and facility scale, interconnection voltages can range from about 13.2 kV for small-scale installations (under 10 MW) to 115-230 kV for hyperscale sites (exceeding 100 MW) in the U.S. [12]. As illustrated in Figure 2, alternating current (AC) power enters the facility through high‑voltage switchgear and protective devices, after which it is stepped down to operational voltage levels such as 480 V via on‑site three‑phase step‑down transformers. These transformers not only reduce voltage but also provide electrical isolation between the utility grid and the facility’s internal power supply system. The power is then distributed through main distribution panels and power distribution units (PDUs) to supply IT server racks, cooling systems, and other supporting loads. For IT server racks, AC power is further converted to direct current (DC) via a power factor correction (PFC) rectifier, followed by a DC-DC converter for DC-side voltage regulation [34], ensuring stable and properly rated power delivery to IT equipment. Additionally, large AI data centers typically employ redundant utility feeders and transformers to ensure high reliability and continued operation even during partial grid outages [35].
@@ II-B2 Uninterruptible Power Supply (UPS)
UPS systems [36] serve as the first layer of protection against power interruptions in data center operations. They provide instantaneous electrical power from energy storage devices, such as batteries or flywheels, when the grid supply is disrupted, ensuring continuous operation during the transition period before backup generators start and synchronize with the load. UPS deployments commonly employ battery technologies such as valve-regulated lead-acid (VRLA), lithium-ion, and emerging solid-state chemistries [37]. Selection criteria include specific power and energy ratings, cycle life, energy density, thermal management requirements, and cost-effectiveness. UPS systems typically operate in three modes: normal mode, double-conversion mode, and battery mode [34], and they are sized to handle high power demands of dense GPU and TPU racks in AI data centers. Modern UPS systems can also support grid-interactive capabilities, enabling data centers to shift loads or participate in demand response programs by temporarily supplying power from their stored energy.
@@ II-B3 Backup Generation
Backup generation provides emergency power to ensure that AI workloads remain uninterrupted in the event of a grid outage. Traditionally, diesel generators have been the most common solution because of low costs and their ability to quickly deliver large amounts of power to reliably support full critical IT load and essential cooling systems. To enhance sustainability, some facilities integrate on-site renewables, such as solar photovoltaic (PV) arrays or hydrogen fuel cell systems, to supplement or partially replace diesel generators [38]. For instance, Caterpillar and Microsoft successfully used hydrogen fuel cells to supply backup power to a data center for 48 hours [39]. These clean generations can operate in parallel with conventional generators during outages and supply power under normal conditions to reduce reliance on the grid. Most backup generation systems are equipped with automatic transfer switches, enabling them to start within seconds and seamlessly supply power during an outage.
@@ II-C Cooling Facilities
In AI data centers, cooling facilities are essential for dissipating the substantial heat generated by high-density computing equipment, particularly GPU- and TPU-based clusters, maintaining safe operation and preventing hardware failures. A variety of cooling technologies have been used in data centers [40], which are broadly classified into three main categories: (i) Air cooling [41] circulates chilled air to absorb heat from IT equipment and expel it from the facility. Due to its low cost and operational simplicity, air cooling has been the mainstream choice for data centers with low rack power densities (below 20 kW), but it becomes inadequate for AI clusters with much higher rack power densities [11]. (ii) Liquid cooling uses a liquid medium to remove heat from servers, which absorbs heat far more effectively than air and allows for very high rack power densities. For example, liquid immersion cooling [42] submerges servers in engineered dielectric fluids, enabling safe, direct contact with electronic components and highly efficient heat removal; direct‑to‑chip liquid cooling [43] circulates coolant through cold plates mounted directly on processors or other high-heat components, improving heat transfer without immersing the entire server in liquid. (iii) Hybrid cooling [44] combines air and liquid cooling methods to optimize efficiency under varying workloads. This approach offers flexible operation, reduces energy consumption, and mitigates the high infrastructure cost and complexity of fully liquid-cooled designs. The commonly used cooling methods are summarized in Table III. In practice, cooling selection depends on factors such as rack power density, energy efficiency targets, facility design constraints, capital and operational costs, and water availability. Particularly, rapidly growing AI workloads are accelerating the adoption of advanced liquid-based and intelligent cooling solutions. Further discussion on smart and efficient cooling techniques is provided in Section V-B.
@@
Air Cooling [41] Low-moderate $\sim$1.1–2.9 [45] Simple, low cost, widely adopted Poor efficiency at high rack power (>>30 kW)
Direct-to-Chip Liquid Cooling [43] High $\sim$1.1–1.3[46] Excellent thermal transfer, ideal for dense AI workloads Requires plumbing and infrastructure; risk of leaks
Immersion Cooling [42] Very high (>>50 kW/rack) $\sim$1.02-1.04 [45] Ultra-efficient, low noise, compact Redesign needed, fluid cost, specialized hardware
Hybrid Cooling [43] Medium-high $\sim$1.34-1.38 [47] Flexible and adaptable to existing setups System integration complexity
@@ III-A Electricity Consumption Structure and Efficiency
AI data centers are emerging as one of the fastest-growing electricity consumers in the energy sector. According to the IEA report [8], global data centers consumed around 415 TWh of electricity in 2024, accounting for about 1.5% of total global electricity consumption. Looking ahead, the IEA projects that global data center electricity demand will more than double by 2030, reaching around 945 TWh, with AI identified as the primary driver of this growth [8]. The global data center electricity consumption is further discussed in Appendix -A.
Modern hyperscale AI facilities typically contract for power capacities exceeding 100 MW, with some campuses planned to scale to the gigawatt level [13], to support ultra-scale AI training and inference. According to the Electric Power Research Institute (EPRI) white paper [14], electricity consumption in a data center is mainly attributed to IT hardware equipment, which typically accounts for 40-50% of total load. Cooling systems represent the second-largest share, consuming approximately 30-40%, depending on the cooling technology and server rack density. The remaining 10-30% is from other supporting facilities, including lighting, office spaces, and monitoring and security systems. In AI data centers, the share of electricity consumed by IT equipment is higher and typically exceeds 60%, due to the adoption of high-density GPU/TPU racks and efficient liquid cooling systems. Power Usage Effectiveness (PUE) [14] is the most widely used metric to measure the energy efficiency of a data center, which is defined as the ratio (1):
@@
A lower PUE indicates greater energy efficiency, with a value of 1.0 signifying that all electricity is used for IT computation. A typical enterprise data center has a PUE of around 1.5-1.6, while modern large-scale facilities achieve a lower PUE below 1.3. For example, Google reports a comprehensive trailing twelve-month PUE of 1.09 across its large-scale data centers, with some sites operating under 1.06 [48]. Similarly, Meta’s 2024 Sustainability Report states an average PUE of 1.08 across its data centers [49]. It indicates that IT computing dominates the electricity consumption of these data centers.

### v4

@@ I Introduction
In recent years, the accelerated advancement of generative artificial intelligence (AI), particularly large language models (LLMs) [zhao2023survey] such as GPT [achiam2023gpt], LLaMA [touvron2023llama], and Gemini [team2023gemini], has fueled explosive growth in the AI industry. This surge is driving the rapid expansion of data center infrastructure and imposing unprecedented pressure on the power grid due to the immense electricity demands of ultra-scale AI workloads. For instance, training GPT-3 is estimated to have consumed 1.29 GWh of electricity [patterson2021carbon], whereas the electricity consumption for training the larger-scale GPT-4 rose dramatically to an estimated over 50 GWh [extnet_gpt4_energy, Cohen2024May], equivalent to nearly 0.1% of New York City’s annual electricity use. According to a recent International Energy Agency (IEA) report [iea2024energyai], global data centers consumed around 415 TWh of electricity in 2024 (about 1.5% of total global demand), and their consumption is projected to more than double by 2030 to around 945 TWh, with AI identified as the primary driver of this growth.
@@
AI data centers are the computing facilities that are designed and optimized to execute large-scale AI workloads, such as the training and inference of LLMs, computer vision systems, and other compute-intensive AI applications. In contrast to traditional data centers, which primarily provide general-purpose information technology (IT) services, AI data centers are architected to deliver the extreme computational performance required for advanced AI workloads (see Table I for a detailed comparison). This capability is enabled by dense configurations of high-performance computing hardware, such as Graphics Processing Units (GPUs) [choquette2021nvidia] and Tensor Processing Units (TPUs) [jouppi2023tpu], together with high-efficiency cooling systems [avelar2023ai]. Modern hyperscale AI data centers typically operate with power demand exceeding 100 MW [semianalysis2024electrical], and some new campuses are planned to scale to the gigawatt (GW) level [crusoe2025wyoming]. As a result, AI data centers are emerging as prominent large electric loads in the power grid, with distinct characteristics and demand patterns, including:
@@
  • High power density: Intensive AI workloads require substantially more power than conventional IT services. For instance, a ChatGPT query is estimated to consume about 2.9 Wh, nearly ten times the 0.3 Wh of a regular Google search [aljbour2024powering]. In contrast to traditional server racks typically operating at 7-10 kW, AI computing racks can reach power densities of 30-100+ kW per rack [nlyte_rack_power]. Such extreme power density places considerable strain on data center architecture, cooling systems, and power grids.
@@
  • Fast and large variability: Across the training, fine-tuning, and inference phases, AI workloads can be highly variable and bursty [Li2024Sep], with power demand fluctuating sharply over very short timescales and remaining difficult to forecast [mughees2025forecast]. For example, large-scale GPU clusters can produce power fluctuations of hundreds of megawatts within only seconds [choukse2025power, aptpower2025ai], bringing significant challenges for system balancing and reliable grid operation.
@@
  • Grid interface via power electronics: AI data centers interface with the power grid primarily through power electronic converters, which exhibit fundamentally different dynamic behaviors from conventional electromechanical loads, such as low inertia, fast response dynamics, and harmonic distortions [Li2025AILoadDynamics]. At scale, power electronics-based AI computing loads can pose major threats to grid stability and cause severe power quality issues [bloomberg2024ai].
@@
  • Geographic concentration: AI data centers tend to cluster in regions with low electricity prices, ample land and water resources, and favorable policy environments. In the U.S., for example, fifteen states (most notably Virginia, Texas, and California) accounted for about 80% of total data center demand in 2023 [aljbour2024powering]. Such concentration effect further intensifies local grid stress, and many regional grids cannot accommodate large-scale integration of data centers without substantial infrastructure upgrades [lin2024exploding].
@@
A growing number of recent articles and reports [Pilz2025Jan, Cohen2024May, iea2024energyai, mckinsey2024datacenters, aljbour2024powering, davenport2024generational] have drawn considerable attention to the surging electricity demand of AI data centers and their implications for the power sector. Prior works [avelar2023ai, aljbour2024powering] examine the energy use of AI data centers by analyzing the power consumption of key components, including AI computing, cooling systems, and auxiliary facilities, and present strategies for energy-efficient operation. Reference [jegham2025hungry] proposes a benchmarking framework for quantifying the footprint of energy, water, and carbon emissions of LLM inference workloads. Recent study [Li2024Sep] analyzes the transient dynamic behaviors of AI computing power consumption, develops high-level mathematical models for AI data center loads, and discusses the resulting disruptions and opportunities for power systems. Case studies on several power grids are conducted in [lin2024exploding] to assess their capacity to accommodate projected five-year AI data center load growth, deriving implications for long-term grid planning and management. However, the existing literature has mainly focused on the AI data center itself and presented high-level, preliminary, and fragmented discussions, while a comprehensive and in-depth synthesis of AI data center load characteristics and their impacts on power grids, particularly the challenges from the grid’s perspective, remains lacking.
@@ II-A1 Computing
AI workloads, such as LLM training and inference, require high-performance computing clusters designed to execute large-scale computations efficiently. These clusters typically integrate multi-core Central Processing Units (CPUs) with dense configurations of GPUs and TPUs [jouppi2023tpu]. These processing units are housed in high-density server nodes and stacked in standardized computer racks, allowing scalable deployment, efficient power distribution, and optimized cooling. Compared with conventional data centers, AI-oriented computing facilities concentrate substantially greater processing power within each rack. While a standard rack typically uses 7-10 kilowatt (kW), an AI-capable rack can demand 30 kW to over 100 kW [aljbour2024powering, davenport2024generational], with an average of more than 60 kW in dedicated AI facilities [nlyte_rack_power]. For example, the per-rack power of NVIDIA’s GB200 NVL36 and NVL72 servers reaches about 66 kW and 120 kW [semianalysisrack], respectively. The substantially higher power densities pose critical challenges for power distribution and thermal management.
@@ II-A2 Storage
To sustain intense compute capabilities, scalable and high‑performance data storage is necessitated to handle massive training datasets, intermediate model checkpoints, and inference data. Tiered storage components include: (i) the hot tier, such as Non-Volatile Memory Express (NVMe) Solid State Drives (SSDs) arrays [do2020cost], provides ultra-low latency and high bandwidth to store active data for fast processing, (ii) the warm tier, such as high-capacity hard disk drives (HDDs), is cost-effective for storing massive, less frequently accessed data, and (iii) the code tier stores data in tape libraries or object storage systems for long-term archiving with less access speed. In addition, distributed file systems, such as Lustre and Ceph [george2025lustre], allow large‑scale parallel reading and writing, while object storage is used for archiving large datasets with long durability and redundancy [AmazonS32024]. Unlike conventional data centers, which balance capacity, cost, and speed, AI-optimized storage prioritizes maximum throughput and parallel access to prevent bottlenecks that idle expensive GPU clusters.
@@ II-A3 Networking
High‑speed networking is critical to ensure that massive data flow efficiently between storage systems and compute nodes during large‑scale AI workloads. The networking system in an AI data center includes both intra-node interconnects, such as NVIDIA NVLink [nvidia_nvlink] for high-bandwidth GPU-to-GPU communication within a server, and inter-node fabrics, such as InfiniBand [infiniband] or Ethernet with RDMA over Converged Ethernet (RoCE), for low-latency, high-throughput communication between servers and racks. Specialized network topologies, including fat‑tree and dragonfly architectures [wilke2020opportunities], are adopted to optimize GPU‑to‑GPU communication in distributed training environments. In contrast, conventional data centers are usually designed for more heterogeneous workloads and general-purpose traffic patterns, and do not require such high bandwidth or low latency.
@@ II-B1 Grid Power Supply
AI data centers are typically connected to the medium- or high-voltage distribution grid, while hyperscale facilities may connect directly to the transmission grid as their primary power source. Depending on the local grid infrastructure and facility scale, interconnection voltages can range from about 13.2 kV for small-scale installations (under 10 MW) to 115-230 kV for hyperscale sites (exceeding 100 MW) in the U.S. [semianalysis2024electrical]. As illustrated in Figure 2, alternating current (AC) power enters the facility through high‑voltage switchgear and protective devices, after which it is stepped down to operational voltage levels such as 480 V via on‑site three‑phase step‑down transformers. These transformers not only reduce voltage but also provide electrical isolation between the utility grid and the facility’s internal power supply system. The power is then distributed through main distribution panels and power distribution units (PDUs) to supply IT server racks, cooling systems, and other supporting loads. For IT server racks, AC power is further converted to direct current (DC) via a power factor correction (PFC) rectifier, followed by a DC-DC converter for DC-side voltage regulation [sun2022dynamic], ensuring stable and properly rated power delivery to IT equipment. Additionally, large AI data centers typically employ redundant utility feeders and transformers to ensure high reliability and continued operation even during partial grid outages [chalise2015data].
@@ II-B2 Uninterruptible Power Supply (UPS)
UPS systems [muhammad2016review] serve as the first layer of protection against power interruptions in data center operations. They provide instantaneous electrical power from energy storage devices, such as batteries or flywheels, when the grid supply is disrupted, ensuring continuous operation during the transition period before backup generators start and synchronize with the load. UPS deployments commonly employ battery technologies such as valve-regulated lead-acid (VRLA), lithium-ion, and emerging solid-state chemistries [teague2025comparative]. Selection criteria include specific power and energy ratings, cycle life, energy density, thermal management requirements, and cost-effectiveness. UPS systems typically operate in three modes: normal mode, double-conversion mode, and battery mode [sun2022dynamic], and they are sized to handle high power demands of dense GPU and TPU racks in AI data centers. Modern UPS systems can also support grid-interactive capabilities, enabling data centers to shift loads or participate in demand response programs by temporarily supplying power from their stored energy.
@@ II-B3 Backup Generation
Backup generation provides emergency power to ensure that AI workloads remain uninterrupted in the event of a grid outage. Traditionally, diesel generators have been the most common solution because of low costs and their ability to quickly deliver large amounts of power to reliably support full critical IT load and essential cooling systems. To enhance sustainability, some facilities integrate on-site renewables, such as solar photovoltaic (PV) arrays or hydrogen fuel cell systems, to supplement or partially replace diesel generators [kambhampati2024moving]. For instance, Caterpillar and Microsoft successfully used hydrogen fuel cells to supply backup power to a data center for 48 hours [gooding2024microsoft]. These clean generations can operate in parallel with conventional generators during outages and supply power under normal conditions to reduce reliance on the grid. Most backup generation systems are equipped with automatic transfer switches, enabling them to start within seconds and seamlessly supply power during an outage.
@@ II-C Cooling Facilities
In AI data centers, cooling facilities are essential for dissipating the substantial heat generated by high-density computing equipment, particularly GPU- and TPU-based clusters, maintaining safe operation and preventing hardware failures. A variety of cooling technologies have been used in data centers [zhang2021survey], which are broadly classified into three main categories: (i) Air cooling [ni2017review] circulates chilled air to absorb heat from IT equipment and expel it from the facility. Due to its low cost and operational simplicity, air cooling has been the mainstream choice for data centers with low rack power densities (below 20 kW), but it becomes inadequate for AI clusters with much higher rack power densities [avelar2023ai]. (ii) Liquid cooling uses a liquid medium to remove heat from servers, which absorbs heat far more effectively than air and allows for very high rack power densities. For example, liquid immersion cooling [kuncoro2019immersion] submerges servers in engineered dielectric fluids, enabling safe, direct contact with electronic components and highly efficient heat removal; direct‑to‑chip liquid cooling [heydari2024experimental] circulates coolant through cold plates mounted directly on processors or other high-heat components, improving heat transfer without immersing the entire server in liquid. (iii) Hybrid cooling [chen2019thermodynamic] combines air and liquid cooling methods to optimize efficiency under varying workloads. This approach offers flexible operation, reduces energy consumption, and mitigates the high infrastructure cost and complexity of fully liquid-cooled designs. The commonly used cooling methods are summarized in Table III. In practice, cooling selection depends on factors such as rack power density, energy efficiency targets, facility design constraints, capital and operational costs, and water availability. Particularly, rapidly growing AI workloads are accelerating the adoption of advanced liquid-based and intelligent cooling solutions. Further discussion on smart and efficient cooling techniques is provided in Section V-B.
@@
Air Cooling [ni2017review] Low-moderate $\sim$1.1–2.9 [haghshenasEnoughHotAir2023] Simple, low cost, widely adopted Poor efficiency at high rack power (>>30 kW)
Direct-to-Chip Liquid Cooling [heydari2024experimental] High $\sim$1.1–1.3[azarifarLiquidCoolingData2024] Excellent thermal transfer, ideal for dense AI workloads Requires plumbing and infrastructure; risk of leaks
Immersion Cooling [kuncoro2019immersion] Very high (>>50 kW/rack) $\sim$1.02-1.04 [haghshenasEnoughHotAir2023] Ultra-efficient, low noise, compact Redesign needed, fluid cost, specialized hardware
Hybrid Cooling [heydari2024experimental] Medium-high $\sim$1.34-1.38 [zhouPowerUsageEffectiveness2022] Flexible and adaptable to existing setups System integration complexity
@@ III-A Electricity Consumption Structure and Efficiency
AI data centers are emerging as one of the fastest-growing electricity consumers in the energy sector. According to the IEA report [iea2024energyai], global data centers consumed around 415 TWh of electricity in 2024, accounting for about 1.5% of total global electricity consumption. Looking ahead, the IEA projects that global data center electricity demand will more than double by 2030, reaching around 945 TWh, with AI identified as the primary driver of this growth [iea2024energyai]. The global data center electricity consumption is further discussed in Appendix -A.
Modern hyperscale AI facilities typically contract for power capacities exceeding 100 MW, with some campuses planned to scale to the gigawatt level [crusoe2025wyoming], to support ultra-scale AI training and inference. According to the Electric Power Research Institute (EPRI) white paper [aljbour2024powering], electricity consumption in a data center is mainly attributed to IT hardware equipment, which typically accounts for 40-50% of total load. Cooling systems represent the second-largest share, consuming approximately 30-40%, depending on the cooling technology and server rack density. The remaining 10-30% is from other supporting facilities, including lighting, office spaces, and monitoring and security systems. In AI data centers, the share of electricity consumed by IT equipment is higher and typically exceeds 60%, due to the adoption of high-density GPU/TPU racks and efficient liquid cooling systems. Power Usage Effectiveness (PUE) [aljbour2024powering] is the most widely used metric to measure the energy efficiency of a data center, which is defined as the ratio (1):
@@
A lower PUE indicates greater energy efficiency, with a value of 1.0 signifying that all electricity is used for IT computation. A typical enterprise data center has a PUE of around 1.5-1.6, while modern large-scale facilities achieve a lower PUE below 1.3. For example, Google reports a comprehensive trailing twelve-month PUE of 1.09 across its large-scale data centers, with some sites operating under 1.06 [googlePUE2025]. Similarly, Meta’s 2024 Sustainability Report states an average PUE of 1.08 across its data centers [metaSustainability2024]. It indicates that IT computing dominates the electricity consumption of these data centers.